Write the equation of the line with the following conditions.
slope = 1/3, passes through (-1,5) in standard form.


We do not have a text book. We are just given worksheets.

Answer by funmath(2873)

You have two formulas you need to know for this:
the point-slope formula: y-y(subscript 1)=m(x-x(subscript 1))
The slope is m and the point is (x(subscript 1),y(subscript 1))
Standard Form: Ax+By=C
in your case m=1/3 and (x(subscript 1), y(subscript 1))=(-1,5)
Plug and chug:
y-5=1/3(x-(-1))
y-5=1/3(x+1)
y-5=1/3(x+1)
y-5=(1/3)x+1/3 clear fractions by multiplying everything by 3.
(3)y+(3)(-5)=(3)(1/3)x+(3)(1/3)
3y-15=x+1
3y-15+15=x+1+15
3y=x+16
-x+3y=-x+x+16
-x+3y=16
Some books don't want the coefficient of x to be negative, so multiply both sides by -1.
-1(-x+3y)=-1(16)
x-3y=-16